I need help with this problem:
2/3e^(4x)+1/3=4

Answer by RAY100(1637)

2/3e^(4x)+1/3=4
subt 1/3 both sides of eqn
2/3e^(4x) =4-1/3=3 2/3=11/3
mult both sides by 3/2
e^(4x)=3/2*11/3= 11/2
ln both sides
4x=ln (11/2)= 1.705
divide both sides by 4
x= .426
check
2/3 e^(4*.426) +1/3=4
2/3 e^(1.705) +1/3 =4
2/3(5.5) +1/3 =4
3 2/3 +1/3 =4
4=4